Grade 9Magical Mathematics[Interesting Approach]

how many zeros will be their in the end of 25!

Shivam Bhagat
15 Years agoGrade 9

1 Answer

AKASH GOYAL AskiitiansExpert-IITD
ApprovedApproved Tutor Answer15 Years ago

Dear Shivam

0 is created by product of 2 and 5

if we can find the exponent of 2 and 5 in the expansion of 25! then it will give number of zero

exponent of prime number 2 in 25!= [25/2] + [25/22]+[25/23]+[25/24]+[25/25]+.......all zero ,  [] denotes greatest integer function here

=12+6+3+1+0... = 22

exponent of 5= [25/5] + [25/52]+[25/53]+0..... = 5+1+0.... = 6

so there are 6 pairs of 2 and 5 it will give rise to 6 zeros in the end

 


All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..